# What is the Relationship between Tensor Factorizations and Circuits (and How Can We Exploit it)?

### v1

@@ 3 From Non-negative Factorizations to Circuits for Probabilistic modeling
Much attention has been posed in machine learning on circuit representations for tractable probabilistic modeling, i.e., for modeling probability distributions that support tractable inference. Circuits built with such a purpose are usually called probabilistic circuits (PCs) (Vergari et al., 2019b; Choi et al., 2020). In this section, we connect non-negative tensor factorizations and PCs, showing a number of research opportunities for the tensor factorization community within the probabilistic machine learning panorama.
@@
Let $p(\bm{\mathrm{X}})$ be a probability mass function (PMF) over finitely-discrete random variables $\bm{\mathrm{X}}=\{X_{j}\}_{j=1}^{d}$, where each $X_{j}\in\bm{\mathrm{X}}$ takes values in $\mathsf{dom}(X_{j})=[I_{j}]$. Then, the simplest representation of $p(\bm{\mathrm{X}})$ is that of a probability tensor $\bm{\mathcal{T}}\in\mathbb{R}_{+}^{I_{1}\times\cdots\times I_{d}}$ such that every entry encodes the probability of a joint configuration of $\bm{\mathrm{X}}$, i.e., $t_{x_{1}\cdots x_{d}}=p(x_{1},\ldots,x_{d})$ for any $\bm{\mathrm{x}}=\langle x_{1},\ldots,x_{d}\rangle\in\mathsf{dom}(\bm{\mathrm{X}})$. Clearly, this representation is inefficient, as it scales exponentially in space with respect to the number of variables $d$. A natural way to compactly model $p(\bm{\mathrm{X}})$ is via a non-negative tensor factorization, e.g., the non-negative version of Tucker (Kim & Choi, 2007), where the factor matrices $\{\bm{\mathrm{V}}^{(j)}\}_{j=1}^{d}$ and the core tensor $\bm{\mathcal{W}}$ showed in Eq. 2 are restricted to have non-negative entries only. By trivially specializing 2, we can encode the non-negative hierarchical Tucker factorization (Vendrow et al., 2021) in a circuit $c$ that outputs non-negative values, also called a PC.
@@
A sufficient condition to ensure a circuit is a PC is constraining both the parameters of sum units and the outputs of input units to be non-negative, resulting in a circuit that is called monotonic (Shpilka & Yehudayoff, 2010). [^8] Non-monotonic PCs – circuits allowing for negative parameters but still guaranteeing non-negative outputs – are possible, see e.g., Loconte et al. (2024a). For instance, the circuit encoding a non-negative hierarchical Tucker factorization that we mentioned above is a monotonic PC, as its sum unit weights (i.e., the entries of the core tensor $\bm{\mathcal{W}}$) and the outputs of its input units (i.e., the entries of the factor matrices $\{\bm{\mathrm{V}}^{(j)}\}_{j=1}^{d}$) are restricted to be non-negative. Smoothness and decomposability in circuits allow for the tractable computation of summation and integrals (Section 2.1), which translates into exactly computing any marginal or conditional distribution for a PC with these structural properties (Vergari et al., 2019b). However, these PCs are not just probabilistic models that allow to tractably evaluate probabilistic queries over $p(\bm{\mathrm{x}})$, they are also generative models from which it is possible to sample exactly.
@@ 3.2 How to parameterize probability tensor factorizations?
Circuits and tensor factorizations are the output of two different optimization problems that however share some common challenges. Understanding them can open new opportunities for both communities. In application scenarios of (non-negative) tensor factorizations, the main task is to compress or reconstruct a given tensor, which is generally explicitly represented in memory. Hence, the parameters of the factorization are optimized as to minimize a reconstruction loss (Cichocki et al., 2007). In contrast, modern PCs are learned from data. That is, one is given a dataset of $N$ datapoints $\{\bm{\mathrm{x}}^{(i)}\}_{i=1}^{N}$ that are assumed to be drawn i.i.d. from and unknown distribution $p(\bm{\mathrm{X}})$ (Bishop & Nasrabadi, 2006). The probability tensor that encodes $p(\bm{\mathrm{X}})$ is therefore implicit and cannot be fully materialized, as the probability distribution is unknown, but also because of its possible exponential—or even infinite—size.
@@
As learning in PCs often reduces to an optimization problem, i.e., maximizing the data (log-)likelihood (Peharz et al., 2016), enforcing the non-negativity of the circuit needs to be done using one or more reparameterizations, i.e., mapping real-valued parameters to the non-negative sum unit weights. This is necessary as the sum unit weights of a monotonic PC need to form a convex combination to yield a valid distribution (as shown in Eq. 8). For instance, we can squash the $K$ parameters $\bm{\theta}\in\mathbb{R}^{K}$ associated to a sum unit with $K$ inputs through a softmax function
$\bm{\mathrm{w}}=\mathsf{softmax}({\bm{\theta}}),\quad\quad\bm{\theta}\in\mathbb{R}^{K}.$ (9)
The usage of such reparameterization combined with input functions encoding probability distributions delivers a PC whose normalization constant is $1$, as the probabilities of all variable assignments sum up to one, which is direct consequence of having the weights of each sum unit summing up to one. For a tensorized circuit, such reparameterization would act row-wise on the parameter matrix associated to every sum layer.
@@
Luckily, if the circuit is smooth and decomposable (Def. 3), we can still compute its normalization constant exactly and efficiently even if sum weights are not normalized (Peharz et al., 2015). This allows us to use alternative ways to reparameterize a monotonic PC $c$, even if its reparameterization delivers an unnormalized distribution, i.e., a distribution not integrating to 1. In fact, we can still recover a valid distribution $p(\bm{\mathrm{X}})$ efficiently via normalization, i.e., $p(\bm{\mathrm{X}})=c(\bm{\mathrm{X}})/Z$ with $Z=\sum_{\bm{\mathrm{x}}\in\mathsf{dom}(\bm{\mathrm{X}})}c(\bm{\mathrm{x}})$ being the normalization constant. For instance, we can enforce each single sum unit parameter $\theta$ to be non-negative via exponentiation
$w=\exp({\theta}),\quad\quad{\theta}\in\mathbb{R}.$ (10)
In this paper, we introduce a third way, a simpler implementation trick that we borrow from the literature on gradient-based optimization for non-negative tensor factorizations (Cichocki et al., 2007): projecting the parameters of all sum units in the positive orthant after every optimization step, i.e.,
$w=\max(\epsilon,\theta),\quad\quad\theta\in\mathbb{R},\ \epsilon>0$ (11)
where $\epsilon$ is a small threshold close to zero and $\max$ is applied element-wise. Each reparameterization can yield a different loss landscape and lead to different solution during optimization. In our experiments (Section 7), we found this third reparameterization to be the most effective to learn monotonic PCs.
@@
When it comes to input units in monotonic PCs, they need to model valid distributions. Common parameterizations can include simple PMFs (or densities, see Section 3.4) such as Bernoulli or Categorical distributions, or even other probabilistic models as long as they can be tractably marginalized. This yields a set of possible parameterizations that go beyond the simple mappings from indices to matrix entries, as usually used in tensor factorizations (1 and 2).
@@
Opportunity 4. A wide range of possible parameterizations Estimating a PMF $p(\bm{\mathrm{X}})$ via a probabilistic model is another way to perform an implicit tensor compression. If this model is a circuit, then this compression exactly maps to a non-negative hierarchical tensor factorization but over a number of basis functions, which are the circuit input units. These input units (thus also input layers in our tensorized formalism) can encode more memory efficient and more expressive functions than indicator functions. For instance, one can use Binomial distributions instead of categoricals as to drastically reduce the number of parameters of the factorizations (Peharz et al., 2020c). In the case of infinite-dimensional probability tensors (see Section 3.4 below), discrete variables with infinite support can instead be modeled by using Poisson distributions (Molina et al., 2017) or generative models as input layers, such as normalizing flows (Papamakarios et al., 2021; Sidheekh et al., 2023), variational auto-encoders (Tan & Peharz, 2019), or also non-linear functions that can be integrated efficiently, e.g., splines (Novikov et al., 2021; Loconte et al., 2024a). Parameterizing input units in this way yields a tensor factorization that uses non-linearities. Along this direction, in the circuit literature parameters of sum layers have been directly parameterized by neural networks (Shao et al., 2020; 2022; Gala et al., 2024a). These non-linear cases have only been explored very recently in the matrix and tensor factorization literature (Leplat et al., 2023; Awari et al., 2024).
@@ 3.3 Reliable Neuro-symbolic integration
A prominent use case for tractable inference with PCs is in safety-critical applications, when it is necessary to enforce hard constraints over the predictions of neural classifiers (Ahmed et al., 2022; van Krieken et al., 2024). These constraints can be expressed as logical formulas over symbols extracted by a perceptual component (the classifier). For example, the safety rule that a self-driving car has to stop in front of a pedestrian or a traffic light (Marconato et al., 2024b; a) can be written as a propositional logical formula $\phi:(\mathrm{P}\vee\mathrm{R}\implies\mathrm{S})$, where $\mathrm{P}$, $\mathrm{R}$ and $\mathrm{S}$ are Boolean variables representing that a $\mathrm{P}$edestrian and a $\mathrm{R}$ed-light have been detected in the video stream of the car and the action to $\mathrm{S}$top has to be taken. Circuits are especially suitable for this neuro-symbolic integration scenario (De Raedt et al., 2019) because they can represent both probability distributions and logical formulas. These two representations can be used in a single classifier to guarantee that the predictions that will violate the given constraint will always have 0 probability. More formally, we can implement such a classifier, mapping inputs $\bm{\mathrm{x}}$ to outputs $\bm{\mathrm{y}}$ that have to satisfy a constraint $\phi$, in this way (Ahmed et al., 2022):
$p(\bm{\mathrm{y}}\mid\bm{\mathrm{x}})\ \propto\ q(\bm{\mathrm{y}}\mid\bm{\mathrm{x}})\mathbbm{1}\!\left\{\bm{\mathrm{y}}\models\phi\right\}$ (12)
@@
Opportunity 5. Structured sparsity via logical constraints Circuits encoding logical formulas are generally very sparse, nonetheless, they still represent a (sparse) factorization of a tensor, in this case a Boolean one. Analogously to the probability tensor described at the beginning of Section 3, this Boolean tensor would encode the logical formula as an exponentially large table of zeros and ones. Multiplying a probability tensor compactly encoded as circuit $q$ as in Eq. 12 with this compact representation of a Boolean tensor equals to a structured form of masking: all the invalid (according to the logical constraint $\phi$) entries in the probability tensor are forcefully set to zero, thus making such entries not predictable. A possible opportunity is therefore to connect with the vast literature of knowledge compilation (Darwiche & Marquis, 2002; Choi et al., 2013; Oztok & Darwiche, 2017) to impose structured sparsity to tensor factorizations. Possible applications include neuro-symbolic integration for graph data (Loconte et al., 2023) as well as representing rankings and user preferences (Choi et al., 2015), scaling cryptographic attacks (Wedenig et al., 2024b), enforcing constraints over the output of LLMs (Zhang et al., 2023) and promoting their self-consistency (Calanzone et al., 2024).
@@ 3.4 Infinite-Dimensional Probability Tensors and Continuous Factorizations
Formally, let $p(\bm{\mathrm{X}})$ be a PDF over continuous variables $\bm{\mathrm{X}}=\{X_{j}\}_{j=1}^{d}$, where each $X_{j}\in\bm{\mathrm{X}}$ takes values in $\mathsf{dom}(X_{j})=\mathbb{R}$. Then, $p(\bm{\mathrm{X}})$ can be represented as an infinite-dimensional probability tensor $\bm{\mathcal{T}}$ such that $t(x_{1},\ldots,x_{d})=p(x_{1},\ldots,x_{d})$ for any $\bm{\mathrm{x}}\in\mathsf{dom}(\bm{\mathrm{X}})$. Infinite-dimensional tensors such as $p(\bm{\mathrm{X}})$ can be decomposed into a finite number of sums and products of factor matrices that live in Hilbert spaces of generic functions. For instance, we can re-adapt the Tucker factorization showed in Def. 1 as a different factorization method where, instead of having factor matrices $\bm{\mathrm{V}}^{(j)}\in\mathbb{R}^{I_{j}\times R_{j}}$ for all $j$, we encode a vector of $R_{j}$ functions $\mathcal{F}^{(j)}=\{f_{r_{j}}^{(j)}\colon\mathsf{dom}(X_{j})\to\mathbb{R}\}_{r_{j}=1}^{R_{j}}$. That is, we factorize $\bm{\mathcal{T}}$ as
@@
Here, we have $\bm{\mathcal{W}}\in\mathbb{R}^{R_{1}\times\cdots\times R_{d}}$. Then, one can trivially modify 1 such that this Tucker factorization of $p(\bm{\mathrm{X}})$ can be represented as a PC of the same size where the input units over variable $X_{j}$ now encode the functions in $\mathcal{F}^{(j)}$. Similarly, one can retrieve PCs encoding mixed probability distributions over discrete and continuous variables (Molina et al., 2018), thus encoding factorizations of a quasi-tensor. In the same way, one can easily re-adapt hierarchical Tucker to factorize $p(\bm{\mathrm{X}})$, thus yielding an equivalent deeper circuits over continuous variables.
Note that, while Eq. 14 is a factorization of an infinite-dimensional tensor, it is still a finite factorization. That is, the ranks $R_{1},\ldots,R_{d}$ are finite, and therefore the circuit representing the same factorization has a sum unit having $R_{1}\cdots R_{d}$ inputs (see Fig. 2). Very recent works have proposed to augment the circuit definition (Def. 2) with integral units which, roughly speaking, encode a sum over an infinite and uncountable number of inputs (Gala et al., 2024a; b). We can consider such PCs as encoding continuous factorizations of a probability tensor, which can be though of as infinite-rank factorizations. For instance, consider the problem of factorizing a finite-dimensional tensor $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$. Instead of considering a finitely-dimensional core tensor $\bm{\mathcal{W}}\in\mathbb{R}^{R_{1}\times\cdots\times R_{d}}$ in Tucker (Eq. 2), we can use a function $\omega\colon\mathsf{dom}(\bm{\mathrm{Z}})\to\mathbb{R}$ over continuous variables $\bm{\mathrm{Z}}=\{Z_{i}\}_{i=1}^{d}$, where each $Z_{i}$ has domain $\mathsf{dom}(Z_{j})=\mathbb{R}$. Similarly, we replace each factor matrix $\bm{\mathrm{V}}^{(j)}\in\mathbb{R}^{I_{j}\times R}$ with a vector of $I_{j}$ functions $\{f_{i_{j}}^{(j)}\colon\mathsf{dom}(Z_{j})\to\mathbb{R}\}_{i_{j}=1}^{I_{j}}$, for all $j$. By doing so and since $\bm{\mathrm{Z}}$ consists of continuous variables, we are in practice replacing the summations in Eq. 2 with a multivariate integral over $\bm{\mathrm{Z}}$. That is, we factorize $\bm{\mathcal{T}}$ as
@@
Similarly, one can retrieve hierarchical versions of such continuous tensor factorizations, with applications for probabilistic modeling (Gala et al., 2024b). In the case of the integral in Eq. 15 being intractable to compute, quadrature rules can be applied as to approximate it. See Gala et al. (2024a) for the details.
@@
In the next section, we present a generic pipeline that can be used to build both finite-dimensional and infinite-dimensional hierarchical probability tensor factorizations as deep tensorized PCs (Def. 7).
4 How to Build and Scale Circuits: A Tensorized Perspective

### v2

@@ 3 From Non-negative Factorizations to Circuits for Probabilistic modeling
Much attention has been paid in machine learning on circuit representations for tractable probabilistic modeling, i.e., for modeling probability distributions that support tractable inference. Circuits built with such a purpose are usually called probabilistic circuits (PCs) (Vergari et al., 2019b; Choi et al., 2020). In this section, we connect non-negative tensor factorizations and PCs, showing a number of research opportunities for the tensor factorization community within the probabilistic machine learning panorama.
@@
Let $p(\bm{\mathrm{X}})$ be a probability mass function (PMF) over finitely-discrete random variables $\bm{\mathrm{X}}=\{X_{j}\}_{j=1}^{d}$, where each $X_{j}\in\bm{\mathrm{X}}$ takes values in $\mathsf{dom}(X_{j})=[I_{j}]$. Then, the simplest representation of $p(\bm{\mathrm{X}})$ is that of a probability tensor $\bm{\mathcal{T}}\in\mathbb{R}_{+}^{I_{1}\times\cdots\times I_{d}}$ such that every entry encodes the probability of a joint configuration of $\bm{\mathrm{X}}$, i.e., $t_{x_{1}\cdots x_{d}}=p(x_{1},\ldots,x_{d})$ for any $\bm{\mathrm{x}}=\langle x_{1},\ldots,x_{d}\rangle\in\mathsf{dom}(\bm{\mathrm{X}})$. Clearly, this representation is inefficient, as it scales exponentially in space with respect to the number of variables $d$. A natural way to compactly model $p(\bm{\mathrm{X}})$ is via a non-negative tensor factorization, e.g., the non-negative version of Tucker (Kim & Choi, 2007), where the factor matrices $\{\bm{\mathrm{V}}^{(j)}\}_{j=1}^{d}$ and the core tensor $\bm{\mathcal{W}}$ shown in Eq. 2 are restricted to have non-negative entries only. By trivially specializing 2, we can encode the non-negative hierarchical Tucker factorization (Vendrow et al., 2021) in a circuit $c$ that outputs non-negative values, also called a PC.
@@
A sufficient condition to ensure a circuit is a PC is constraining both the parameters of sum units and the outputs of input units to be non-negative, resulting in a circuit that is called monotonic (Shpilka & Yehudayoff, 2010).333 Non-monotonic PCs, which allow negative weights while ensuring non-negative outputs, are possible Loconte et al. (2024). For instance, the circuit encoding a non-negative hierarchical Tucker factorization that we mentioned above is a monotonic PC, as its sum unit weights (i.e., the entries of the core tensor $\bm{\mathcal{W}}$) and the outputs of its input units (i.e., the entries of the factor matrices $\{\bm{\mathrm{V}}^{(j)}\}_{j=1}^{d}$) are restricted to be non-negative. Smoothness and decomposability in circuits allow for the tractable computation of summation and integrals (Section 2.1), which translates into exactly computing any marginal or conditional distribution for a PC with these structural properties (Vergari et al., 2019b). However, these PCs are not just tractable probabilistic models, they are also generative models from which it is possible to sample exactly.
@@ 3.2 How to Parameterize Probability Tensor Factorizations?
Circuits and tensor factorizations are the output of two different optimization problems that however share some common challenges. Understanding them can open new opportunities for both communities. In application scenarios of (non-negative) tensor factorizations, the main task is to compress or reconstruct a given tensor, which is generally explicitly represented in memory. Hence, the parameters of the factorization are optimized as to minimize a reconstruction loss (Cichocki et al., 2007). In contrast, modern PCs are learned from data. That is, one is given a dataset of $N$ datapoints $\{\bm{\mathrm{x}}^{(i)}\}_{i=1}^{N}$ that are assumed to be drawn i.i.d. from and unknown distribution $p(\bm{\mathrm{X}})$ (Bishop & Nasrabadi, 2006). The probability tensor that encodes $p(\bm{\mathrm{X}})$ is therefore implicit and cannot be fully materialized, as the probability distribution is unknown, but also because of its possible exponential size (or even infinite, see Section 3.4).
@@
As learning in PCs often reduces to an optimization problem, i.e., maximizing the data (log-)likelihood (Peharz et al., 2016), enforcing the non-negativity of the circuit is done by using one or more reparameterizations, i.e., mapping real-valued parameters to positive sum unit weights. This is necessary as the sum weights of a monotonic PC need to form a convex combination to yield a valid distribution (as shown in Eq. 8). For instance, we can squash the $K$ parameters $\bm{\theta}\in\mathbb{R}^{K}$ of a sum unit with $K$ inputs through a softmax function, i.e.  $\bm{\mathrm{w}}=\mathsf{softmax}({\bm{\theta}})$. Using such a reparameterization together with input functions encoding probability distributions delivers a PC whose normalization constant is $1$, as the probabilities of all variable assignments sum up to one. This is direct consequence of having the weights of each sum unit summing up to one. For tensorized circuits, this reparameterization would act row-wise on the parameter matrix of every sum layer.
@@
Luckily, if the circuit is smooth and decomposable (Def. 3), we can still compute its normalization constant exactly and efficiently even if sum weights are not normalized (Peharz et al., 2015). This allows us to use alternative ways to reparameterize a monotonic PC $c$, even if its reparameterization delivers an unnormalized distribution, i.e., a distribution not integrating to 1. In fact, we can still recover a distribution $p(\bm{\mathrm{X}})$ efficiently via normalization, i.e., $p(\bm{\mathrm{X}})=c(\bm{\mathrm{X}})/Z$ with $Z=\sum_{\bm{\mathrm{x}}\in\mathsf{dom}(\bm{\mathrm{X}})}c(\bm{\mathrm{x}})$ being the normalization constant. For instance, we can enforce each sum unit parameter $\theta$ to be non-negative via exponentiation, i.e. $w=\exp({\theta})$. In this paper, we introduce a third way, a simpler implementation trick that we borrow from the literature on gradient-based optimization for non-negative tensor factorizations (Cichocki et al., 2007): projecting the sum unit parameters in the positive orthant after every optimization step, i.e.,
$w=\max(\epsilon,\theta),\quad\quad\theta\in\mathbb{R}$ (9)
@@
where $\epsilon$ is a positive threshold close to zero. Each reparameterization can yield a different loss landscape and lead to different solution during optimization. In our experiments (Section 6), we found this third reparameterization to be the most effective to learn PCs. When it comes to input units in monotonic PCs, they need to model valid distributions. Common parameterizations can include simple PMFs (or densities, see Section 3.4) such as Bernoulli or Categorical distributions, or even other probabilistic models as long as they can be tractably marginalized. This yields a set of possible parameterizations that go beyond the simple mappings from indices to matrix entries, as usually used in tensor factorizations (1 and 2).
@@
Opportunity 4. A wide range of possible parameterizations Estimating a PMF $p(\bm{\mathrm{X}})$ via a probabilistic model is another way to perform an implicit tensor compression. If this model is a circuit, then this compression exactly maps to a non-negative hierarchical tensor factorization but over a number of basis functions, which are the circuit input units. These input units (thus also input layers in our tensorized formalism) can encode more memory efficient and more expressive functions than indicators. For instance, one can use Binomial distributions instead of categoricals as to drastically reduce the number of parameters of the factorizations (Peharz et al., 2020c). In the case of infinite-dimensional probability tensors (see Section 3.4 below), discrete variables with infinite support can instead be modeled by using Poisson distributions (Molina et al., 2017) or generative models as input layers, such as normalizing flows (Papamakarios et al., 2021; Sidheekh et al., 2023), variational auto-encoders (Tan & Peharz, 2019), or also non-linear functions that can be integrated efficiently, e.g., splines (Novikov et al., 2021; Loconte et al., 2024). Parameterizing input units in this way yields a tensor factorization that uses non-linearities. Along this direction, in the circuit literature parameters of sum layers have been directly parameterized by neural networks (Shao et al., 2020; 2022; Gala et al., 2024a). These non-linear cases have only been explored very recently in the matrix and tensor factorization literature (Leplat et al., 2023; Awari et al., 2024).
@@ 3.3 Reliable Neuro-Symbolic Integration
A prominent use case for tractable inference with PCs is in safety-critical applications, where it is necessary to enforce hard constraints over the predictions of neural classifiers (Ahmed et al., 2022; van Krieken et al., 2024). Such constraints can be expressed as logical formulas over symbols extracted by a perceptual component (a classifier). For example, the safety rule that a self-driving car must stop in front of a pedestrian or a traffic light (Marconato et al., 2024b; a) can be written as a propositional logical formula $\phi:(\mathrm{P}\vee\mathrm{R}\implies\mathrm{S})$, where $\mathrm{P}$, $\mathrm{R}$ and $\mathrm{S}$ are Boolean variables representing that a $\mathrm{P}$edestrian and a $\mathrm{R}$ed-light have been detected in the video stream of the car and the action to $\mathrm{S}$top must be taken.
Circuits are especially suitable for this neuro-symbolic integration (De Raedt et al., 2019), because they can represent both probability distributions and logical formulas. These two representations can be used in a single classifier to guarantee that the predictions that will violate the given constraint will always have 0 probability. Formally, we can implement such a classifier, mapping inputs $\bm{\mathrm{x}}$ to outputs $\bm{\mathrm{y}}$ that have to satisfy a constraint $\phi$, as (Ahmed et al., 2022):
$p(\bm{\mathrm{y}}\mid\bm{\mathrm{x}})\ \propto\ q(\bm{\mathrm{y}}\mid\bm{\mathrm{x}})\mathbbm{1}\!\left\{\bm{\mathrm{y}}\models\phi\right\},$ (10)
@@
Opportunity 5. Structured sparsity via logical constraints Circuits encoding logical formulas are generally very sparse, nonetheless, they still represent a (sparse) factorization of a tensor, in this case a Boolean one. Analogously to the probability tensor described at the beginning of Section 3, this Boolean tensor would encode the logical formula as an exponentially large table of zeros and ones. Multiplying a probability tensor compactly encoded as circuit $q$ as in Eq. 10 with this compact representation of a Boolean tensor equals to a structured form of masking: all the invalid (according to the logical constraint $\phi$) entries in the probability tensor are forcefully set to zero, thus making such entries not predictable. A possible opportunity is therefore to connect with the vast literature of knowledge compilation (Darwiche & Marquis, 2002; Choi et al., 2013; Oztok & Darwiche, 2017) to impose structured sparsity to tensor factorizations. Possible applications include neuro-symbolic integration for graph data (Loconte et al., 2023) as well as representing rankings and user preferences (Choi et al., 2015), scaling cryptographic attacks (Wedenig et al., 2024b), enforcing constraints over the output of LLMs (Zhang et al., 2023) and promoting their self-consistency (Calanzone et al., 2025).
@@ 3.4 Infinite-Dimensional Probability Tensors and Continuous Factorizations
Formally, let $p(\bm{\mathrm{X}})$ be a PDF over continuous variables $\bm{\mathrm{X}}=\{X_{j}\}_{j=1}^{d}$, where each $X_{j}\in\bm{\mathrm{X}}$ takes values in $\mathsf{dom}(X_{j})=\mathbb{R}$. Then, $p(\bm{\mathrm{X}})$ can be represented as an infinite-dimensional probability tensor $\bm{\mathcal{T}}$ such that $t(x_{1},\ldots,x_{d})=p(x_{1},\ldots,x_{d})$ for any $\bm{\mathrm{x}}\in\mathsf{dom}(\bm{\mathrm{X}})$. Infinite-dimensional tensors such as $p(\bm{\mathrm{X}})$ can be decomposed into a finite number of sums and products of factor matrices that live in Hilbert spaces of generic functions. For instance, we can re-adapt the Tucker factorization shown in Def. 1 as a different factorization method where, instead of having factor matrices $\bm{\mathrm{V}}^{(j)}\in\mathbb{R}^{I_{j}\times R_{j}}$ for all $j$, we encode a vector of $R_{j}$ functions $\mathcal{F}^{(j)}=\{f_{r_{j}}^{(j)}\colon\mathsf{dom}(X_{j})\to\mathbb{R}\}_{r_{j}=1}^{R_{j}}$. That is, we factorize $\bm{\mathcal{T}}$ as
@@
Here, we have $\bm{\mathcal{W}}\in\mathbb{R}^{R_{1}\times\cdots\times R_{d}}$. Then, one can trivially modify 1 such that this Tucker factorization of $p(\bm{\mathrm{X}})$ can be represented as a PC of the same size where the input units over variable $X_{j}$ now encode the functions in $\mathcal{F}^{(j)}$. Similarly, one can retrieve PCs encoding mixed probability distributions over discrete and continuous variables (Molina et al., 2018), thus encoding factorizations of a quasi-tensor. In the same way, one can easily re-adapt hierarchical Tucker to factorize $p(\bm{\mathrm{X}})$, thus yielding an equivalent deep circuit over continuous variables.
Note that, while Eq. 12 is a factorization of an infinite-dimensional tensor, it is still a finite factorization. That is, the ranks $R_{1},\ldots,R_{d}$ are finite, and therefore the circuit representing the same factorization has a sum unit having $R_{1}\cdots R_{d}$ inputs (see Fig. 2). Very recent works have proposed to augment the circuit definition (Def. 2) with integral units which, roughly speaking, encode a sum over an infinite and uncountable number of inputs (Gala et al., 2024a; b). We can consider such PCs to encode continuous factorizations of a probability tensor, which can be though of as infinite-rank factorizations. For instance, consider the problem of factorizing a finite-dimensional tensor $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$. Instead of considering a finitely-dimensional core tensor $\bm{\mathcal{W}}\in\mathbb{R}^{R_{1}\times\cdots\times R_{d}}$ in Tucker (Eq. 2), we can use a function $\omega\colon\mathsf{dom}(\bm{\mathrm{Z}})\to\mathbb{R}$ over continuous variables $\bm{\mathrm{Z}}=\{Z_{i}\}_{i=1}^{d}$, where each $Z_{i}$ has domain $\mathsf{dom}(Z_{j})=\mathbb{R}$. Similarly, we replace each factor matrix $\bm{\mathrm{V}}^{(j)}\in\mathbb{R}^{I_{j}\times R}$ with a vector of $I_{j}$ functions $\{f_{i_{j}}^{(j)}\colon\mathsf{dom}(Z_{j})\to\mathbb{R}\}_{i_{j}=1}^{I_{j}}$, for all $j$. By doing so and since $\bm{\mathrm{Z}}$ consists of continuous variables, we are in practice replacing the summations in Eq. 2 with a multivariate integral over $\bm{\mathrm{Z}}$. That is, we factorize $\bm{\mathcal{T}}$ as
@@
Similarly, one can retrieve hierarchical versions of such continuous tensor factorizations, with applications for probabilistic modeling (Gala et al., 2024b). In case the integral in Eq. 13 is intractable to compute, quadrature rules can be applied as to approximate it. See Gala et al. (2024a) for the details.
In the following section (Section 4), we present a generic pipeline that can be used to build both finite-dimensional and infinite-dimensional hierarchical probability tensor factorizations as deep tensorized PCs (Def. 7). Before that, in the following opportunity box, we stress how circuits can also be used as alternative representations of probability distributions that do not correspond to probability tensor factorizations.